# Understanding Retrieval Robustness for Retrieval-Augmented Image Captioning

### v1

@@ Abstract
Recent advancements in retrieval-augmented models for image captioning highlight the significance of retrieving related captions for efficient, lightweight models with strong domain-transfer capabilities. While these models demonstrate the success of retrieval augmentation, retrieval models are still far from perfect in practice. Retrieved information can sometimes mislead the model generation, negatively impacting performance. In this paper, we analyze the robustness of the SmallCap retrieval-augmented captioning model. Our analysis shows that SmallCap is sensitive to tokens that appear in the majority of the retrieved captions, and integrated gradients attribution shows that those tokens are likely copied into the final caption. Given these findings, we propose to train the model by sampling retrieved captions from more diverse sets. This reduces the probability that the model learns to copy majority tokens and improves both in-domain and cross-domain performance effectively.
@@ 1 Introduction
Recent retrieval-augmented image captioning models have shown success in strong image captioning performance while reducing model parameters by retrieving related captions for a given image (Ramos et al., 2023b; Sarto et al., 2022; Yang et al., 2023). These models use retrieved information as additional context besides the input image. However, similar to retrieval-augmented language models (Yoran et al., 2023), image captioning models enhanced with retrieval can sometimes be misled by irrelevant information. For example, in Figure 1 the captioning model is distracted by the word “elephant” in the retrieved captions and generates captions that do not match the given image.
For retrieval-augmented language models, Yoran et al. (2023) have studied the misled results from retrieved context, and propose to train a retrieval-robust LLM by continuous training with synthetic data for question answering tasks. However, in their approach, the retriever returns only one passage at each step. Considering that LLMs can be sensitive to the order of prompts Lu et al. (2022), the robustness of using multiple retrieved results has not been fully studied. Evaluating and improving the robustness of retrieval-augmented image captioning models remains under-explored, specifically when the model is augmented with multiple retrieved results.
Figure 1: Comparison of generated image captions that are predicted without retrieval, misled by retrieval, and predicted with a more retrieval-robust model.
To bridge this gap in the literature, we study the robustness of a retrieval-augmented captioning model. By the definition of retrieval robustness proposed in Yoran et al. (2023), retrieved context should boost model performance when relevant, and should not adversely affect it when irrelevant. We thoroughly examine the model robustness regarding the order of the retrieved captions, and the relevance of the retrieved content. We further propose a majority voting analysis incorporating attribution and attention to investigate how the retrieved tokens impact the model generation. And finally, inspired by Hoang et al. (2022), we propose to sample the retrieved captions from a larger list during training to prevent the model from overfitting to the top relevant captions. Our evaluation shows improved model robustness and better out-of-domain generalization.
@@ 2 Related Work
Recent efforts aim to enhance RAG model robustness against misguided or hallucinated generations. One approach involves filtering retrieved content (Wang et al., 2023; Yoran et al., 2023; Yasunaga et al., 2023; Yan et al., 2024; Asai et al., 2023) by applying or training an additional evaluator. Another direction focuses on improving robustness during the training of the generation model itself. Specifically, for retrieval-augmented question answering with large language models, Yoran et al. (2023) propose continued training with a mixed synthetic dataset, while Cuconasu et al. (2024) suggests incorporating irrelevant documents. In retrieval-augmented translation, robustness is improved through shuffling retrieved translations (Hoang et al., 2022), ensemble model decoding (Hao et al., 2023), and controlled interactions between source and retrieved translations (Hoang et al., 2023).
@@
Image captioning is the task that describes the visual contents of an image in natural language Xu et al. (2015); Osman et al. (2023). Recent studies have integrated RAG into this field. Sarto et al. (2022) and Zhou and Long (2023) experimented with retrieving similar or style-aware images before generating captions. Li et al. (2023) introduced a lightweight image captioning model that utilizes retrieved concepts. More related to our work, Ramos et al. (2023a) developed end-to-end encoder-decoder models that simultaneously attend to both the image and retrieved caption embeddings.
@@ 3 Robustness of Retrieval-Augmented Image Captioning
To evaluate the robustness of the existing state-of-the-art retrieval-augmented caption model, SmallCap Ramos et al. (2023b), we conduct controlled experiments and observe its resilience to changes in (1) the order of the retrieved captions and (2) the content relevance of the retrieved captions.
Figure 2: CIDEr evaluation of smallcap on the COCO validation set when using the top-$k$, low(er)-ranked, randomly retrieved captions, against a baseline without retrieval augmentation. Performance drops by up to 50% when using randomly retrieved captions compared to no-rag, suggesting that that the model is not robust.
@@ 3.1 Robustness Evaluation
For a given image, SmallCap is augmented with a sequence of $k$ retrieved captions that are provided into a prompt for the decoder: “Similar images show $cap_{1},cap_{2},cap_{3},cap_{k}$. This image shows …”. The retrieved captions are obtained through image-to-text retrieval using CLIP embeddings, and are ordered according to their relevance, i.e., cosine similarity. From the sorted retrieved captions, we retain the most similar captions as the retrieval list. In this regard, the top-$k$ relevant candidates are the first $k$ captions in the list, and the low-ranked captions are the last-captions in the list. SmallCap uses the top-$k$ relevant captions in the prompt.
@@
When prompting the model to generate a caption for a given image, we change the order of the retrieved captions by permuting or reversing them. We evaluate the effect of the order changes in two settings: one with a model trained using the top-$k$ relevant captions (default), and another that is also trained with permuted or reversed retrieved captions. [^2] For the model trained with default order—top four captions, we use the pretrained checkpoints from HuggingFace: https://huggingface.co/Yova/SmallCap7M and https://huggingface.co/Yova/SmallCapOPT7M
@@
To evaluate how robust the model is towards noise in the retrieved captions, we are curious to see how the model performs when (1) captions are randomly retrieved, i.e. the retrieved captions are irrelevant to the given image (2) only low-ranked retrieved captions are available. Here the randomly retrieved captions are those retrieved with another image. For low-ranked captions, we take the lowest-ranked $k$ captions from the retrieval list that consists of top seven relevant captions.
@@ 3.2 Experimental Setup
In this experiments, we set $k=4$ as it has been demonstrated as the optimal number of captions by Ramos et al. (2023b). We evaluate SmallCap models with both OPT-350m Zhang et al. (2022) and GPT-2 Radford et al. (2019) as the decoder models. For the image encoder, we use ResNet-50x64 He et al. (2016) and CLIP-ViT-B/32 Radford et al. (2021) as the retrieval encoder. We keep the same model setting in the following sections unless stated otherwise.
@@
We first evaluate robustness of Smallcap on COCO validation set for In-domain evaluation. Then we evaluate on NoCaps (Agrawal et al., 2019), which contains In, Near and Out-domain samples, and serves as a challenging dataset designed to assess the generalization capabilities of models trained on COCO. For both datasets we use the validation set experimenting with different number of retrieved captions, i.e. different $k$ values. We report CIDEr (Vedantam et al., 2015).
Retrieval Order SmallCap Variant
Training Evaluation GPT-2 OPT
@@
Table 1: CIDEr score when manipulating the order of the top-k retrieved captions by either randomly permuting or reversing the list. We evaluate on the COCO validation set using either a GPT-2 or OPT backbone in the SmallCap model.
@@ 3.3 Order Robust but Content Sensitive
From Table 1 and Table 2 we observes that, smallcap is relatively robust to the order of the retrieved texts. Permuting the order of the retrieved captions during training and evaluation show $1$ CIDEr point improvement for COCO (Lin et al., 2014) and $2-3$ CIDEr score increase for Nocaps (Agrawal et al., 2019). This indicates that if multiple retrieved results are used as augmentation, permuting their order helps.
@@
As shown in Figure 2, when using randomly retrieved captions instead of the top-$k$ relevant ones during inference, performance drops drastically compared to the no retrieval baseline. [^3] Here the top and low ranked captions are obtained from a list of top seven retrieved captions that are ordered by their cosine similarity to the image embedding. This implies that smallcap lacks resilience to noise in the retrieved captions, and the irrelevant context has the potential to misguide the model, resulting in inaccurate caption predictions. When prompting with low-ranked retrieved captions, while performance slightly decreases, the retrieval-augmented model still outperforms the one without retrieval.
SmallCap Variant
Retrieval Order GPT-2 OPT
@@
Table 2: CIDEr score on the NoCaps dataset when trained and evaluated with different order of the top-four retrieved captions.
4 Majority Tokens Explain Behaviour
To better understand how each token of the retrieved content relates to the observed sensitivity discussed in the previous section, we hypothesize that the model predictions are driven by the presence of majority tokens. In other words, when the captioning model is prompted with retrieved captions, we assume the predicted tokens are influenced by the tokens that appear in the majority of the retrieved captions. To verify the assumption, we designed a majority voting analysis, followed by attribution and attention analysis for model behavior interpretation.
@@ 4.1 Majority Tokens
We first introduce the definition of majority tokens. Let $R=[T_{1},\cdots,T_{n}]$ represent the retrieved caption $R$, which contains a list of tokens. Assuming we retrieved $N$ captions for a given image, for each token $T_{i}$ in captions $R_{1}$, $R_{2}$, …, $R_{N}$, we define the token $T_{i}$ as a majority token (denoted as $T_{M}$) if $T_{i}$ appears in more than half of the retrieved captions, i.e. $C_{T_{i}}>\frac{N}{2}$ where $C_{T_{i}}$ is the frequency of the token appearance as in Equation 1:
$C_{T_{i}}=\sum_{n=1}^{N}\mathbbm{1}_{T_{i}\in R_{n}}$ (1)
For each predicted caption $Y$ of the evaluation samples, we can calculate the majority-vote probability $P_{T_{M}\in Y}$ as the probability of the majority token $T_{M}$ existing in the tokens of the predicted caption $Y$, where $Y=[Y_{1},\cdots,Y_{n}]$.
It is intuitive that the higher the value of $P_{T_{M}\in Y}$, the more likely the model is generating captions with a majority voting mechanism within the text context, as the model is relying more on the overlap of the tokens in the retrieved captions.
@@ 4.2 Experimental Setup
To verify the majority vote assumption, we design a controlled experiment. Specifically, we analyze the predictions of the model when it is provided with three retrieved captions (i.e., $N=3$) where
  • setup 2G1B contains two relevant captions and one irrelevant caption;
  • setup 2B1G contains two irrelevant captions and one relevant caption.
The assumption is that, if there is a majority voting mechanism among the tokens in the retrieved captions, the model will copy such majority tokens to the final prediction. The distinction will be obvious in this setting — in the setup 2B1G, if the model is robust to the retrieved context, the model will focus more on the relevant retrieval instead of the majority tokens, as the majority tokens root in the two irrelevant captions.
We use the COCO evaluation set and the pretrained checkpoint with the OPT decoder of Ramos et al. (2023b) for this experiment. The retrieved captions are obtained following Ramos et al. (2023b). For the controlled setups, we utilize the top one and top two captions retrieved from the COCO captions data store for the given image, respectively. The irrelevant captions are obtained by adding the retrieved captions from another randomly selected image. Note that we remove the stop words in the retrieved captions when collecting majority tokens. The stop words are filtered from the top-100 most frequent tokens in the COCO dataset, where we manually remove meaning tokens such as “man”, “two” from the list. Please see the Appendix A for the complete list.
@@ Results.
By evaluating the pretrained model on the defined setups, we found that the probability of majority vote is 86.47% for setup 2G1B and 20.84% for setup 2B1G. This indicates that the majority tokens in the relevant captions are strong hints for model generation. For the majority tokens that exist in the irrelevant captions, the model has some robustness to not always follow them. However, in 20.84% of the time it is misled by their appearance and such majority tokens is therefore copied to the generated captions. In other words, the majority vote probability gives us an upper bound error estimation for the model when the retrieved results are noisy and misleading.
@@ 4.3 Input Attribution with Integrated Gradients
We use integrated gradients (Sundararajan et al., 2017) for input attribution analysis to determine the role of majority tokens in the model predictions. This enables us to examine the influence of each individual token in the retrieved captions on the final prediction during the generation process.
@@
Figure 3 shows an attribution visualization where the attribution score of each input token (x-axis) is computed at each generation step (y-axis). Bright color cells correspond to high attribution to the input token. High attribution scores to the same tokens seen in the retrieved captions indicates probability of copying. Negative attribution scores are observed at contradicting tokens observed in the retrieved captions to the current generation. For example, negative scores are observed at token “her” when model is predicting the token “boy” and at token “small” when predicting “herd”. Additional input attribution visualisations can be found in the Appendix.
@@
We also quantitatively analyze the impact of majority tokens by calculating pairwise attribution scores between tokens in retrieved captions and those predicted by the model. Higher attribution values suggest greater sensitivity to the input token Ancona et al. (2018). Figure 4 shows the distribution of the pairwise attribution scores for the 2B1G setup. It is clear that the model is sensitive to the majority tokens, especially when the generated token can be copied from the retrieval set. Such behavior indicates weak robustness: we would not expect a robust model to be distracted by the tokens from the two irrelevant retrieved sentences at inference time. To better visualize the impact, we show distribution of original attribution values and the absolute values Ancona et al. (2018) across all evaluation samples.
Figure 4: Pairwise average attribution score between retrieved and generated tokens. MT: majority tokens in the retrieved captions in the 2B1G setup. OT: other tokens. The larger pairwise attribution values shows that the majority tokens have a larger impact during generation than the other tokens in the retrieved captions.

### v3

@@ Abstract
Recent advances in retrieval-augmented models for image captioning highlight the benefit of retrieving related captions for efficient, lightweight models with strong domain-transfer capabilities. While these models demonstrate the success of retrieval augmentation, retrieval models are still far from perfect in practice: the retrieved information can sometimes mislead the model, resulting in incorrect generation and worse performance. In this paper, we analyze the robustness of a retrieval-augmented captioning model SmallCap. Our analysis shows that the model is sensitive to tokens that appear in the majority of the retrieved captions, and the input attribution shows that those tokens are likely copied into the generated output. Given these findings, we propose to train the model by sampling retrieved captions from more diverse sets. This decreases the chance that the model learns to copy majority tokens, and improves both in-domain and cross-domain performance.
@@ 1 Introduction
Recent retrieval-augmented image captioning models have shown success in strong image captioning performance while reducing model parameters by retrieving related captions for a given image (Ramos et al., 2023b; Sarto et al., 2022; Yang et al., 2023). These models use retrieved information as additional context besides the input image. However, similar to retrieval-augmented language models (Yoran et al., 2023), image captioning models enhanced with retrieval can sometimes be misled by irrelevant information. For example, in Figure 1 the captioning model is misled by the token “elephant” in the retrieved captions, and generates captions that do not match the given image.
For retrieval-augmented language models, Yoran et al. (2023) have studied the cases where retrieval misled the model prediction, and address this problem with a retrieval-robust LLM by continuous training with synthetic data for question answering tasks. However, in their approach, the retrieval system returns only one passage at each step. Considering that LLMs can be sensitive to the order of prompts Lu et al. (2022), the robustness of using multiple retrieved results has not been fully studied. Evaluating and improving the robustness of retrieval-augmented image captioning models remains under-explored, specifically when the model is augmented with multiple retrieved results.
Figure 1: Comparison of generated image captions that are predicted without retrieval, misled by retrieval, and predicted with a more retrieval-robust model. The retrieval-augmented model generates the token “elephant”, which appears in 3/4 of the retrieved captions.
To bridge this gap in the literature, we study the robustness of the SmallCap retrieval-augmented captioning model (Ramos et al., 2023b). By the definition of retrieval robustness proposed in Yoran et al. (2023), retrieved context should boost model performance when relevant, and should not adversely affect it when irrelevant. We thoroughly examine the robustness of the model with regards to the order of the retrieved captions, and the relevance of the retrieved content. We also present a novel analysis of model behaviour based on majority voting, supported by input attribution and attention analyses to investigate how the retrieved tokens influence the model generation. And finally, inspired by Hoang et al. (2022), we propose to sample the retrieved captions from a larger list during training to prevent the model from overfitting to the top relevant captions. Our evaluation shows improved model robustness and better out-of-domain generalization.
@@ 2 Related Work
Recent efforts aim to enhance RAG model robustness against misguided or hallucinated generations. One approach involves filtering retrieved content (Wang et al., 2023; Yoran et al., 2023; Yasunaga et al., 2023; Yan et al., 2024; Asai et al., 2023) by applying or training an additional evaluator. Another direction focuses on improving robustness during the training of the generation model itself. Specifically, for retrieval-augmented question answering with large language models, Yoran et al. (2023) propose continued training with a synthetic dataset that contains both relevant and irrelevant context, while Cuconasu et al. (2024) suggests incorporating irrelevant documents. In retrieval-augmented translation, robustness is improved through shuffling retrieved translations (Hoang et al., 2022), ensemble model decoding (Hao et al., 2023), and controlled interactions between source and retrieved translations (Hoang et al., 2023).
@@
Image captioning is the task that describes the visual contents of an image in natural language Xu et al. (2015); Osman et al. (2023). Recent studies have integrated RAG into this field. Sarto et al. (2022) and Zhou and Long (2023) experimented with retrieving similar or style-aware images before generating captions. Li et al. (2023) introduced a lightweight image captioning model that utilizes retrieved concepts. More related to our work, Ramos et al. (2023a) developed end-to-end encoder-decoder models that attend to both the image and retrieved caption embeddings.
@@ 3 Robustness of Retrieval-Augmented Image Captioning
To evaluate the robustness of the SmallCap retrieval-augmented caption model Ramos et al. (2023b), we conduct controlled experiments and observe its resilience to changes in (1) the order of the retrieved captions and (2) the content relevance of the retrieved captions.
Figure 2: CIDEr evaluation of SmallCap on the COCO validation set using the top-$k$, low(er)-ranked, randomly retrieved captions, against a baseline without retrieval augmentation. Performance drops by up to 50% when using randomly retrieved captions compared to baseline, suggesting that the model is not robust.
@@ 3.1 Robustness Evaluation
For a given image, SmallCap is augmented with a sequence of $k$ retrieved captions that are combined into an input for the language model decoder: “Similar images show $cap_{1},cap_{2},...,cap_{k}$. This image shows …”. The retrieved captions are obtained through image-to-text retrieval using CLIP embeddings (Radford et al., 2021), and are sorted according to their relevance, i.e., cosine similarity. From the sorted retrieved captions, we retain the most similar captions as the retrieval list. In this regard, the top-$k$ retrieved candidates are the first $k$ captions in the list, and the low-ranked captions are the last-$k$ captions in the list. SmallCap uses the top-$k$ retrieved captions in the prompt by default.
@@
When prompting the model to generate a caption for a given image, we can change the order of the retrieved captions by permuting or reversing them. We evaluate the effect of the order changes in two settings: one with a model trained using the top-$k$ retrieved captions (default), and another that is also trained with permuted or reversed retrieved captions. [^2] For the model trained with default order—top four captions, we use the pretrained checkpoints from HuggingFace: https://huggingface.co/Yova/SmallCap7M, https://huggingface.co/Yova/SmallCapOPT7M
@@
To evaluate how robust the model is towards noise in the retrieved captions, we are curious to see how the model performs when (1) captions are randomly retrieved, i.e. likely to be irrelevant for the given image (2) only low-ranked retrieved captions are available. Here the randomly retrieved captions are those retrieved with another image. For low-ranked captions, we take the lowest-ranked $k$ captions from the retrieval list that consists of top seven relevant captions.
@@ 3.2 Experimental Setup
In the experiments, we set $k=4$ as it has been demonstrated as the optimal number of captions by Ramos et al. (2023b). We evaluate SmallCap models with both OPT-350M Zhang et al. (2022) and GPT-2 Radford et al. (2019) as the decoder models. For the image encoder, we use ResNet-50x64 He et al. (2016) and CLIP-ViT-B/32 Radford et al. (2021) as the retrieval encoder. We keep the same model setting in the following sections unless stated otherwise.
@@
We first evaluate the robustness of Smallcap on COCO validation set for in-domain evaluation. Then we evaluate on NoCaps (Agrawal et al., 2019), which contains In, Near and Out-of-domain data, and serves as a challenging dataset designed to assess the generalization capabilities of models trained on COCO. For both datasets we use the validation set experimenting with different number of retrieved captions, i.e. different $k$ values. We report peformance using CIDEr score (Vedantam et al., 2015).
Retrieval Order LM Backbone
Train Eval GPT-2 OPT
@@
Table 1: CIDEr evaluation on the COCO validation set with GPT-2 and OPT variants of SmallCap when manipulating the order of the top-k retrieved captions.
@@ 3.3 Order Robust but Content Sensitive
From the results in Table 1 and Table 2, we observe that SmallCap is indeed robust to the order of the retrieved texts. Permuting the order of the captions during training and evaluation show $1$ CIDEr point improvement for COCO (Lin et al., 2014) and $2-3$ CIDEr score increase for NoCaps (Agrawal et al., 2019). This indicates that if multiple captions are used for augmentation, then permuting their order helps.
@@
Figure 2 shows that when using randomly retrieved captions instead of the top-$k$ most relevant captions, performance drops drastically compared to the no-retrieval baseline. [^3] Here the top and low ranked captions are obtained from a list of top-seven captions retrieved captions ordered by their cosine similarity to the image embedding. This implies that smallcap lacks resilience to noise in the retrieved captions, and the irrelevant context has the potential to mislead the model, resulting in inaccurate predictions. When prompting with low-ranked retrieved captions, while performance slightly decreases, the retrieval-augmented model still outperforms the one without retrieval.
Retrieval Order LM Backbone
GPT-2 OPT
@@
Table 2: Evaluation on NoCaps using CIDEr score with the GPT-2 and OPT variants of SmallCap when manipulating the order of the top-$k$ retrieved captions.
4 Majority Tokens Explain Behavior
To better understand how each token of the retrieved content relates to the observed sensitivity discussed in the previous section, we hypothesize that the model is driven by the presence of majority tokens. In other words, when the model is prompted with retrieved captions, we assume that the predicted tokens are influenced by the tokens that appear in the majority of the retrieved captions. To test this assumption, we propose a majority voting analysis, followed by input attribution, and an attention analysis of the model behavior.
@@ 4.1 Majority Tokens
We first introduce the definition of majority tokens. Let $R=[T_{1},\cdots,T_{n}]$ represent a retrieved caption $R$, which contains a sequence of $n$ tokens. For a given image, we assume that a total of $K$ retrieved captions are used in the model prompt: $R_{1}$, $R_{2}$, …, $R_{K}$. For each token $T_{i}$ in the set of unique tokens from the retrieved captions, we define $T_{i}$ as a majority token (denoted as $T_{M}$) if $T_{i}$ appears in more than half of the retrieved captions [^4] Note that we remove the stop words in the retrieved captions when determining the majority tokens. The stop words are filtered from the top-100 most frequent tokens in the COCO dataset, where we manually remove meaningful tokens such as “man”, “two” from the list. Please see the Appendix A for the complete list., i.e., $C_{T_{i}}>\frac{K}{2}$ where $C_{T_{i}}$ is the number of retrieved captions that contains token ${T_{i}}$ as in Equation 1:
$C_{T_{i}}=\sum_{l=1}^{K}\mathbbm{1}[T_{i}\in R_{k}]$ (1)
For a generated caption $Y=[y_{1},\cdots,y_{n}]$ in the evaluation data, we can calculate the majority-vote probability $P_{T_{M}\in Y}$ as the probability of the majority token $T_{M}$ appearing in the generated caption.
We expect that the higher the value of $P_{T_{M}\in Y}$, the more likely it is that the model is generating captions based on the majority tokens.
@@ 4.2 Experimental Setup
We test our majority vote assumption with a controlled experiment. Specifically, we analyze the predictions of the model in two settings, each provided with $K=3$ retrieved captions to ensure the presence of a majority token:
2 Good 1 Bad (2G1B):
The retrieval set contains two relevant captions and one irrelevant caption;
2 Bad 1 Good (2B1G):
The retrieval set contains two irrelevant captions and one relevant caption.
The assumption is that, if there is a majority voting behavior with respect to the retrieved captions, the model will copy such majority tokens to the final output. The distinction will be clear in this setting — in the setup 2B1G, if the model is robust to the retrieved context, the model will focus more on the good caption instead of the majority tokens in the two bad captions.
We use the COCO evaluation set and the pretrained checkpoint with the OPT decoder of Ramos et al. (2023b) for this analysis. Good captions are obtained using the top-two and top-one retrieved captions, respectively, for a given image. Bad captions are obtained by retrieving one or two captions, respectively, from a randomly selected image.
@@ Results.
We find that the probability of majority vote in the 2G1B setting is 86.47%. This high probability suggests that the majority tokens in the good captions could be being used to guide the model generation. In the 2B1G setting, the model is much less likely to generate majority tokens from the bad captions, indicating some robustness in not always following them. However, 20.84% of the time, the model can still be misled by their appearance, resulting in the majority tokens being copied into the model output.
@@ 4.3 Input Attribution with Integrated Gradients
To better understand the role of majority tokens in model generation, we use integrated gradients (Sundararajan et al., 2017) for input attribution analysis. This enables us to examine the influence of each individual token in the retrieved captions on the model prediction.
@@
Figure 3 shows an example of an attribution visualization, where the attribution score of each input token (x-axis) is computed at each generation step (y-axis). Bright color cells correspond to high attribution to the input token. High attribution scores to the same tokens seen in the retrieved captions may indicate copying. Negative attribution scores are observed at contradicting tokens observed in the retrieved captions to the current generation. Negative scores are observed at token “her” when model is predicting the token “boy” and at token “small” when predicting “herd”. Additional input attribution visualizations can be found in Appendix B.1.
@@
We also quantitatively analyze the impact of majority tokens by calculating pairwise attribution scores between tokens in retrieved captions and those predicted by the model. Higher attribution values suggest greater sensitivity to the input token Ancona et al. (2018). Figure 4 shows the distribution of the pairwise attribution scores for the 2B1G setup. It is clear that the model is sensitive to the majority tokens, especially when the generated token exists in the retrieved captions. Such behavior indicates weak robustness: we would not expect a robust model to be distracted by the tokens from the two irrelevant retrieved sentences at inference time. To better visualize the impact, we show distribution of original attribution values and the absolute values Ancona et al. (2018) across all evaluation samples.
Figure 4: Pairwise average attribution score between retrieved and generated tokens in the 2B1G setup. MT: majority tokens in the retrieved captions. OT: all other tokens. The larger pairwise attribution values shows that the majority tokens have a larger impact during generation than the other tokens in the retrieved captions.